OpenStudy (anonymous):

Max wants to make an acrylic paper weight shaped like a right square pyramid, as shown below. What is the volume of acrylic required to make the paper weight?

OpenStudy (anonymous):

giving me the error again... wait a few more seconds....

OpenStudy (anonymous):

ok.. got it....

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

the volume of any pyrmid = (1/3)*(area of base)*(height) here, h= 12 what's the area of base?

OpenStudy (anonymous):

49?

OpenStudy (anonymous):

yes... plug it in...

OpenStudy (anonymous):

49 times 12=588/3=196

OpenStudy (anonymous):

correct...:)

OpenStudy (anonymous):

the answer is 196cm cubed

OpenStudy (anonymous):

oh yeah... don't forget the units....

OpenStudy (anonymous):

k, next one?

OpenStudy (anonymous):

sure... close up....:)